OpenStudy (anonymous):

Solve for x then calculate xy: 3x+4y=23 4x-3y=14 a. xy= -6 b. xy= 10 c. xy= 9 d. xy= 24 e. none of the above

OpenStudy (anonymous):

how fast you need it?

OpenStudy (anonymous):

i'm getting C for my answer. @satellite73 i'm just trying to do some practice work as extra study for an upcoming exam. whenever is okay.

OpenStudy (shadowlegendx):

If you need it fast, @satellite73 is the perfect dewd for you ;3

OpenStudy (shadowlegendx):

Though there have been rumors that @satellite73 is a girl c;

OpenStudy (anonymous):

for once it is NOT C (amazing)

OpenStudy (anonymous):

\[3x+4y=23\\ 4x-3y=14\] multiply the first equation 3, the second by 4 and the y terms will drop out \[9x+12y=69\\ 16x-12y=56\] add and get \[25x=125\]making \(x=5\)

OpenStudy (shadowlegendx):

Interesting concept @satellite73